OpenStudy (ellie202000):

Perform the indicated operation 4/7x^2-3/2x^3 I got 1/14x^2

OpenStudy (anonymous):

4/49−3/2x^3

OpenStudy (ellie202000):

How?

OpenStudy (anonymous):

use PEMDAS, Orders of Operations and Simplify

OpenStudy (ellie202000):

Well my options are\[A) \frac{ -8x+21 }{ 14x ^{3}}\]\[B) \frac{ 8x+21 }{ 14x ^{3} }\]\[C) \frac{ -8+21 }{ 14x ^{3} }\]\[D) \frac{ 8x-21 }{ 14x ^{3} }\]

OpenStudy (ellie202000):

@SolomonZelman

OpenStudy (ellie202000):

or anyone

OpenStudy (ellie202000):

@campbell_st

OpenStudy (campbell_st):

well you need a common denominator... so given 7x^2 and 2x^3 the common denominator will be 14x^3 so that's the main task of the question \[\frac{4}{7x^2} \times \frac{2x}{2x} - \frac{3}{2x^3} \times \frac{7}{7} = \frac{8x}{14x^3} - \frac{21}{14x^3}\] I'll let write it in a simpler form hope it helps

OpenStudy (ellie202000):

Yes it does. Thank you